I need help with simplifying these radical expressions please! I don't know how to put the sign on here, but its radical 128, I need to know how to simplify it. Also, radical 48a^2b^3c^4..Thanks Alot
Answer by RAY100(1637)

sqrt(128)=sqrt(2*64)=sqrt(64)*sqrt(2)=+/- 8sqrt2
remember factoring radicals is searching for perfect squares that factor into expression
sqrt (48a^2b^3c^4)=sqrt ( (16*3)a^2 b^2 b^1 c^4 = +/- 4abc^2 sqrt (3b) answer
remember sqrt b^3 = sqrt (b^2 *b^1)=sqrt b^2 * sqrt b^1= (b)sqrt(b)